OpenStudy (anonymous):

Steps on how to establish this identity: (tanθ+secθ-1)/(tanθ-secθ+1) = tanθ+secθ

OpenStudy (aravindg):

let me do on paper first

OpenStudy (aravindg):

good question ...Gimme some more time ..Do you have any ideas?

OpenStudy (anonymous):

That's fine! I'm still working on it... I tried rewriting the left side and I ended up with (sinθ-cosθ+1)/sineθ+cosθ-1), but I'm stuck

OpenStudy (aravindg):

ya I first thought of that ..But it seems there is no need for that both LHS and RHS have tan theta and cos theta ..So we dont need to write tan theta in terms of sin and cos

OpenStudy (aravindg):

*tan theta and sec theta

OpenStudy (aravindg):

I think I got it :)

OpenStudy (anonymous):

YES! :D

OpenStudy (aravindg):

u also gt it ? :)

OpenStudy (anonymous):

Maybe, I'm still checking though

OpenStudy (aravindg):

ok first you say ur answer then I will say mine

OpenStudy (anonymous):

Ok, it's going to be really long, so you'll have to wait a while as I type lol

OpenStudy (aravindg):

yep I will :)

OpenStudy (aravindg):

...........

OpenStudy (aravindg):

lol

OpenStudy (anonymous):

\[\frac{\tanθ+(\secθ-1)}{\tanθ- (\secθ-1)}· \frac{\tanθ+(\secθ-1)}{\tanθ+(\secθ-1)}= \frac{ \tan²θ+2\tanθ(\secθ-1)+\sec²θ-2\secθ+1}{\tan²θ-(\sec²θ-2\secθ+1)}\]

OpenStudy (anonymous):

wth that took forever

OpenStudy (aravindg):

wow :) I did exactly the same

OpenStudy (aravindg):

nah not really ..hw abt this

OpenStudy (aravindg):

What I did \[\dfrac{\tan \theta +(\sec \theta -1)}{\tan \theta -(\sec \theta -1) }=\dfrac{\tan \theta +(\sec \theta -1)}{\tan \theta -(\sec \theta -1) } \times \dfrac{\tan \theta +(\sec \theta -1)}{\tan \theta +(\sec \theta -1) }\] \[=\dfrac{\tan^2 \theta +\sec ^2 \theta +1-2\sec \theta +2\tan \theta(\sec \theta -1)}{\tan^2 \theta-\sec^2 \theta +2\sec \theta -1}\] \[=\dfrac{2\sec^2 \theta-2 \sec \theta +2\tan \theta(\sec \theta-1)}{2(\sec \theta-1)}\] \[=\dfrac{2\sec \theta(\sec \theta-1)+2\tan \theta(\sec \theta-1)}{2(\sec \theta-1)}\] Cancelling off \(2(\sec \theta-1)\) on numerator and denominator . =\(\tan \theta +\sec \theta=RHS \)

OpenStudy (anonymous):

WOW yours looks amazing! :) Better than mine! I was like okay first step and you're done lol. Typing in the equations took forever.. it amazes me how you could type that!

OpenStudy (aravindg):

;)

OpenStudy (aravindg):

Anyway glad that you figured it out yourself!

OpenStudy (anonymous):

It took a while! :) Thanks so much! I appreciate your help!

OpenStudy (aravindg):

yw :)